Circular No. 8184

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN CRUX V. Tabur, Wanniassa, A.C.T., reports his discovery on CCD images

(taken with a 140-mm f/2.8 Nikon telephoto lens) of a possible nova of mag 10.2 at α = 12^{h}23^{m}16^{s}.2, δ = −60^{o}22′34′′ (equinox 2000.0; estimated uncertainty ± 4′′) on Aug. 20.486 and 20.490 UT. The object was possibly visible near limiting mag 11.9 on Aug. 18, but not visible on Tabur’s many CCD images back to Jan. 2000. A brief observation through clouds by Tabur around Aug. 21.5 indicated that the object had brightened by ≈ 1 mag in one day. L. A. G. Monard, Pretoria, S. Africa, reports that the new object appeared at unfiltered CCD mag 9.2 on Aug. 21.71. Both observers add that nothing appears at this position on Palomar survey plates.

SUPERNOVAE 2003hg, 2003hh, 2003hi, 2003hj, 2003hk, 2003hl Several additional apparent supernovae have been reported from un-

filtered CCD images — 2003hg, 2003hh, 2003hk, and 2003hl by M. Moore and W. Li (via LOTOSS/KAIT; cf. IAUC 8182), and 2003hi, 2003hj, and 2003hk by T. Boles (cf. IAUC 8179). The tabulated data below for 2003hg, 2003hh, and 2003hl are from Li, the rest from Boles.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003hg Aug. 18.4 23^{h}51^{m}24.^{s}13 +20^{o}06′38′′.3 16.9 11′′.5 W, 3′′.9 S
2003hh Aug. 18.5  0 00 07.08 + 8 16 42.3 18.7 0′′.4 E, 2′′.9 S
2003hi Aug. 19.87 15 56 33.80 +41 53 47.5 18.1 8′′.9 E, 6′′.1 S
2003hj Aug. 19.92 15 24 20.51 +29 57 11.9 17.5 1′′.0 E, 8′′.2 S
2003hk Aug. 20.11 2 46 25.69 + 3 36 32.3 17.6 5′′.8 E, 6′′.0 N
2003hl Aug. 20.5  1 59 21.28 +19 00 14.5 16.5 24′′.2 E, 13′′.0 S

Additional approximate magnitudes (from KAIT unless denoted ‘B’ for Boles): SN 2003hg in NGC 7771, Aug. 9.4 UT, [19.0; 19.4, 16.9. SN 2003hh in UGC 12890, Aug. 8.5, [19.5; 19.4, 18.1. SN 2003hi in MCG +07-33-16, Feb. 28, [19.0 (B); Apr. 8, [19.5 (B); May 28, [19.5 (B); Aug. 20.852, 18.1 (B). SN 2003hj in MCG +05-36-28, Feb. 3, [19.0 (B); Apr. 8, [19.5 (B); May 28, [19.5 (B); Aug. 20.846, 17.9 (B). SN 2003hk in NGC 1085, Jan. 7, [19.0 (B); Feb. 1.2, [19.0; 17, [19.0 (B); Aug. 20.5, 16.1; 21.063, 17.6 (B); 21.5, 16.0. SN 2003hl in NGC 772, Aug. 11.5, [19.0; 21.5, 16.4. SN 2003hi is also absent from Palomar Sky Survey 1991 red and 1988 blue plates, SN 2003hj from 1987 red and 1990 blue plates, and SN 2003hk from 1990 red and 1992 blue plates. Position end figures for SN 2003hk from Li: 25^{s}.76, 32′′.2 (or 6′′.3 east and 5′′.9 north of the nucleus of NGC 1085).

2003 August 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green